### v1

@@ Abstract
Facial affective behavior analysis (FABA) is crucial for understanding human mental states from images. However, traditional approaches primarily deploy models to discriminate among discrete emotion categories, and lack the fine granularity and reasoning capability for complex facial behaviors. The advent of Multi-modal Large Language Models (MLLMs) has been proven successful in general visual understanding tasks. However, directly harnessing MLLMs for FABA is challenging due to the scarcity of datasets and benchmarks, neglecting facial prior knowledge, and low training efficiency. To address these challenges, we introduce (i) an instruction-following dataset for two FABA tasks, i.e., facial emotion and action unit recognition, (ii) a benchmark FABA-Bench with a new metric considering both recognition and generation ability, and (iii) a new MLLM EmoLA as a strong baseline to the community. Our initiative on the dataset and benchmarks reveal the nature and rationale of facial affective behaviors, i.e., fine-grained facial movement, interpretability, and reasoning. Moreover, to build an effective and efficient FABA MLLM, we introduce a facial prior expert module with face structure knowledge and a low-rank adaptation module into pre-trained MLLM. We conduct extensive experiments on FABA-Bench and four commonly-used FABA datasets. The results demonstrate that the proposed facial prior expert can boost the performance and EmoLA achieves the best results on our FABA-Bench. On commonly-used FABA datasets, EmoLA is competitive rivaling task-specific state-of-the-art models. The dataset and codes will be released: https://anonymous.4open.science/r/FABA-EmoLA-6C76/.
@@ 2.2 Multi-modal LLMs and efficient LLM adaptation
Multi-modal LLMs. Multi-modal LLMs are getting popular in multi-modal content understanding [11, 117, 23, 111, 64]. They are built on top of LLMs [16, 108, 14, 133, 2, 42], and transform visual (videos and images) and text data into a sequence of tokens as input, resulting in generative modeling of downstream multi-modal understanding tasks by next token prediction. Specific to the image-based MLLMs, image tokens are typically encoded by CLIP vision transformer [89, 24]. Then, one of the major challenges in MLLM is how to project image token features into the language domain to better utilize the instruction-following capability of LLMs. In literature, Flamingo [1] is an early work that uses cross-attention to build the interaction between images and text tokens. This design is followed by recent MLLMs [49, 101, 3, 13]. Based on the cross-attention mechanism, Q-Former [51] was proved to be a superior visual-text projector and inspires recent line of MLLM research [145, 7, 19]. Recently, in contrast to the attention-based projector, LLaVA series [68, 67] propose to use a simple MLP as the projector and resort to the LLMs to handle the visual-text interaction. In this paper, based on LLaVA-1.5 [67], we developed an MLLM for affective behavior analysis and benchmarked multiple MLLM baselines.
Efficient LLM adaptation. Though most existing multimodal LLMs show impressive performance on general visual-language tasks, how to adapt them to downstream applications is still challenging [120], especially with limited instructional annotation data and computing resources. This raises the surge of research on parameter-efficient fine-tuning (PEFT) [37], which avoids fine-tuning the entire LLM in training. According to [120], existing PEFT methods can be categorized into addictive [37, 65, 34, 91, 87, 76, 48, 57, 66], partial [128, 32, 121, 103], reparametrized [38, 21, 130, 70], hybrid [34, 44, 77, 9], and unified fine-tuning [113, 36]. In this paper, we are interested in the re-parameterized fine-tuning as it gains increasing research attention. In this line, LoRA [38] is a pioneering work, which is inspired by the fact that modal weight adaptation has a low intrinsic rank. In training, LoRA keeps most LLM parameters frozen while only optimizing the low-rank factorized matrices of dense layers’ residual. In our work, we empirically found the effectiveness of LoRA on the adaptation of MLLMs toward affective behavior analysis tasks.
@@ 3 Dataset and Benchmark
Table 2: FABA-Instruct statistics. Statistics Value Total images 19877 Emotion training samples 19474 Emotion testing samples 403 Emotion description average length 50.47 AU training samples 15835 AU testing samples 350 AU description average length 207.35 Refer to caption Figure 2: FABA-Instruct annotation.
@@ 3.1 Instruction-following FABA dataset
Action Unit description analysis. The annotations from traditional AUR datasets, e.g., BP4D [136], DISFA [79], EmotioNet [27], GFT [29], mainly adopt a string of binary vector to denote whether each AU is activated or not. However, the representation capability of this way is limited. It cannot indicate the degree of AU’s activation and cannot provide any explanations and analysis for the prediction. For instance, in Fig. 4, merely stating that AU6 (cheek raiser) is activated cannot depict the degree of cheek muscle raising. In contrast, the AU descriptions can provide more reasoning cues like “small squint” or “crow’s feet”, which can capture the activated degree of AU6. Furthermore, descriptions can also show the inference ability by providing the relationship between current AU and other AUs or emotions, e.g., “indicative of a frown or a scowl”, “associated with strong emotions”. Such descriptions can not only provide more nuances but also improve the interpretability of the model on AU tasks.
@@ 3.2 Instruction-following FABA benchmark
Our REGE score is defined to consider both text generation and image recognition aspects of FABA models. For text generation, we choose to use the ROUGE score, which is a generally used metric in NLG by comparing the overlap of n-grams, word sequences, and word pairs between the generated texts and the reference ones. For recognition, we adopt the recognition accuracy for the multi-class performance of facial emotion recognition, and the F1 score for the multi-label performance of AU recognition. Denote the recognition performance as $S_{re}$ and the generation metric as $S_{ge}$, our REGE metric is computed by taking their sum: $S_{rege}=S_{ge}+S_{re}$.
@@ 4 EmoLA: An Instruction-tuned MLLM for FABA
After obtaining the visual embedding tokens $H_{v}$, facial prior token $H_{p}$ and the language embedding tokens $H_{q}$, we concatenate them together and feed them into the LLM decoder. Here we use Vicuna [14] as the LLM decoder. As shown in Fig. 7, the visual expert, prior encoder, word embedding and the LLM decoder are frozen. Except for the prior encoder $f_{p}(\cdot)$, the weights of the other modules come from a pretrained MLLM. In this paper, we utilize the LLaVA-1.5 [67] as the backbone model. In order to train efficiently without tuning the entire LLM, we propose to tune an extra LoRA module $h_{\phi}(\cdot)$ and the prior projector $g_{\theta}(\cdot)$. Compared to finetuning the entire LLM, tuning LoRA reduces both memory and computation costs during training without inducing additional expenses for inference. Our experiments demonstrate the effectiveness of such a design in Tab. 7. As a result, the overall parameters to be optimized are $\boldsymbol{\Theta}=\{\theta,\phi\}$. We optimize these parameters following the auto-regressive way, and the likelihood of the generated description $X_{A}$ conditioned on image $X_{V}$, prior facial feature $Z_{P}$ and instructions $X_{Q}$ is given by:
$\vspace{-5pt}p\left(X_{A}|X_{V},Z_{P},X_{Q}\right)=\prod_{i=1}^{L}{p_{\boldsymbol{\Theta}}\left(x_{i}|X_{V},Z_{P},X_{Q},X_{A,<i}\right)},$ (2)
@@ 5.2 Comparison on FABA-Bench
We compare with current MLLMs, i.e., LLaVA-1.5 [67], MiniGPT4-V2 [10], Shikra [11], and mPLUG-Owl2 [125], on our FABA-Bench in Tab. 7. These baselines mainly employ Vicuna [14] as LLM decoder and incorporate instruction tuning. We reproduce all the baselines using their open-sourced codes. For fair comparison, all the models are trained for 1 epoch and train on two tasks individually. Except for our EmoLA, all the other baselines are finetuned based on the pretrained MLLMs. More details about these baselines can refer to Appendix.
It is noteworthy that EmoLA achieves the best results in two tasks, and specifically outperforms other MLLMs by a significant margin on AUR task. EmoLA finetunes merely 10% of the parameters compared to LLaVA-1.5, yet it achieves better performance on FABA-Bench, which can be attributed to two aspects. For one thing, EmoLA only finetunes LoRA which is much more efficient than tuning the entire LLM decoder. For another, EmoLA incorporates a facial prior expert to extract the facial structure knowledge. This feature compensates for the information overlooked by the visual encoder in FABA tasks. Moreover, we can also see that due to the strong capability of LLM, the language generation
Table 8: Influence of prior token. $H_{v}$ and $H_{p}$ indicate the visual tokens and facial prior token, respectively.
@@
64.5 31.7 60.2 34.7
Table 9: Influence of tuning strategy. $h_{\phi}$ and $h_{\theta}$ indicate the LoRA module and the prior projector, respectively.
@@
64.5 31.7 60.2 34.7
ability of Shikra, LLaVA-1.5 and EmoLA is comparable, making it unreasonable to evaluate these models merely based on the NLG metrics. Our metric, REGE, accounts for both the recognition and generation abilities of MLLMs in FABA, offering a more comprehensive evaluation of their performance.
5.3 Ablation study
Compared to merely inputting the prior token $H_{p}$, incorporating visual tokens $H_{v}$ will significantly enhance recognition performance. These visual tokens carry more detailed face information, which will benefit FABA tasks. Incorporating both visual tokens and prior tokens leads to further improvements with eligible performance dropping in text generation. This suggests that the face prior token supplies complementary information of visual tokens. We suppose this additional information includes facial structure details introduced by landmark facial prior extractor.
Figure 8: An example of EmoLA’s capability on FABA tasks.
We also tried exclusively tuning LoRA $h_{\phi}$ while keeping other parameters frozen. This is different from merely inputting the visual tokens $H_{v}$ since it also includes a prior token $H_{p}$ generated by the randomly initialized prior projector $g_{\theta}$. The results indicate that tuning $h_{\phi}$ significantly improves the performance by better aligning the output of the LLM. It can also be observed that the performance slightly exceeds that achieved using only the visual tokens, which also emphasizes the value of the prior token. Again, the efficacy is further augmented by simultaneously fine-tuning both LoRA and the prior projector. As a result, we adopt tuning both two modules during the training process.
@@ 0.A.1 Limitations
This work also has its limitations. Firstly, we haven’t tried other face-prior feature extractors except for landmark features. We leave this as an exploring research direction for future work. Secondly, our annotations from the training set also include noise induced by GPT-4V’s hallucinations [105, 106] (see Sec. 0.C.3), which may introduce some bias for models. We think our benchmark can be regarded as the learning from noisy descriptions task.
@@ 0.C.3 The accuracy estimation of training annotations
Table 11: Estimated accuracy of Emotion annotations in FABA-Instruct. Total Correct Acc. 200 182 91.0 Table 12: Estimated F1 score of AU annotations in FABA-Instruct.
AU1 AU2 AU4 AU5 AU6 AU7 AU9 AU10 AU12 AU15 92.2 95.6 84.2 90.7 82.6 65.9 61.3 58.1 84.3 70.6
AU17 AU20 AU23 AU24 AU25 AU26 AU27 AU43 Avg. 66.7 55.2 79.1 79.5 84.9 83.9 78.8 57.1 76.1
@@ 0.D.2 Baseline details of FABA-Bench
EmoLA (single task) 64.5 31.7 96.2 60.2 34.7 94.9
@@ 0.D.3 Multi-task performance on FABA-Bench
EmoLA (prior token before visual tokens) 63.5 32.1 95.6 57.8 35.3 93.1
EmoLA (prior token after visual tokens) 64.5 31.7 96.2 60.2 34.7 94.9
@@ 0.E.2 Failed cases
(a) Examples of EmoLA’s failed prediction on emotion task.

### v2

@@ Abstract
Facial affective behavior analysis (FABA) is crucial for understanding human mental states from images. However, traditional approaches primarily deploy models to discriminate among discrete emotion categories, and lack the fine granularity and reasoning capability for complex facial behaviors. The advent of Multi-modal Large Language Models (MLLMs) has been proven successful in general visual understanding tasks. However, directly harnessing MLLMs for FABA is challenging due to the scarcity of datasets and benchmarks, neglecting facial prior knowledge, and low training efficiency. To address these challenges, we introduce (i) an instruction-following dataset for two FABA tasks, i.e., facial emotion and action unit recognition, (ii) a benchmark FABA-Bench with a new metric considering both recognition and generation ability, and (iii) a new MLLM EmoLA as a strong baseline to the community. Our initiative on the dataset and benchmarks reveal the nature and rationale of facial affective behaviors, i.e., fine-grained facial movement, interpretability, and reasoning. Moreover, to build an effective and efficient FABA MLLM, we introduce a facial prior expert module with face structure knowledge and a low-rank adaptation module into pre-trained MLLM. We conduct extensive experiments on FABA-Bench and four commonly-used FABA datasets. The results demonstrate that the proposed facial prior expert can boost the performance and EmoLA achieves the best results on our FABA-Bench. On commonly-used FABA datasets, EmoLA is competitive rivaling task-specific state-of-the-art models. The dataset and codes are available: https://johnx69.github.io/FABA/.
@@ 2.2 Multi-modal LLMs and efficient LLM adaptation
Multi-modal LLMs. Multi-modal LLMs are getting popular in multi-modal content understanding [11, 123, 24, 117, 67]. They are built on top of LLMs [17, 114, 15, 140, 2, 43], and transform visual (videos and images) and text data into a sequence of tokens as input, resulting in generative modeling of downstream multi-modal understanding tasks by next token prediction. Specific to the image-based MLLMs, image tokens are typically encoded by CLIP vision transformer [94, 25]. Then, one of the major challenges in MLLM is how to project image token features into the language domain to better utilize the instruction-following capability of LLMs. In literature, Flamingo [1] is an early work that uses cross-attention to build the interaction between images and text tokens. This design is followed by recent MLLMs [50, 107, 3, 14]. Based on the cross-attention mechanism, Q-Former [52] was proved to be a superior visual-text projector and inspires recent line of MLLM research [152, 7, 20, 59]. Recently, in contrast to the attention-based projector, LLaVA series [71, 70] propose to use a simple MLP as the projector and resort to the LLMs to handle the visual-text interaction. In this paper, based on LLaVA-1.5 [70], we developed an MLLM for affective behavior analysis and benchmarked multiple MLLM baselines.
Efficient LLM adaptation. Though most existing multimodal LLMs show impressive performance on general visual-language tasks, how to adapt them to downstream applications is still challenging [126], especially with limited instructional annotation data and computing resources. This raises the surge of research on parameter-efficient fine-tuning (PEFT) [38], which avoids fine-tuning the entire LLM in training. According to [126], existing PEFT methods can be categorized into addictive [38, 68, 35, 96, 92, 80, 49, 58, 69], partial [135, 33, 127, 109], reparametrized [39, 22, 137, 73], hybrid [35, 45, 82, 9], and unified fine-tuning [119, 37, 134, 103]. In this paper, we are interested in the re-parameterized fine-tuning as it gains increasing research attention. In this line, LoRA [39] is a pioneering work, which is inspired by the fact that modal weight adaptation has a low intrinsic rank. In training, LoRA keeps most LLM parameters frozen while only optimizing the low-rank factorized matrices of dense layers’ residual. In our work, we empirically found the effectiveness of LoRA on the adaptation of MLLMs toward affective behavior analysis tasks.
@@ 3 Dataset and Benchmark
Table 2: FABA-Instruct statistics.
Statistics Value
Total images 19877
Emotion training samples 19474
Emotion testing samples 403
Emotion description average length 50.47
AU training samples 15838
AU testing samples 325
AU description average length 207.35
Figure 2: FABA-Instruct annotation.
@@ 3.1 Instruction-following FABA dataset
Action Unit description analysis. The annotations from traditional AUR datasets, e.g., BP4D [143], DISFA [84], EmotioNet [28], GFT [30], mainly adopt a string of binary vector to denote whether each AU is activated or not. However, the representation capability of this way is limited. It cannot indicate the degree of AU’s activation and cannot provide any explanations and analysis for the prediction. For instance, in Fig. 4, merely stating that AU6 (cheek raiser) is activated cannot depict the degree of cheek muscle raising. In contrast, the AU descriptions can provide more reasoning cues like “small squint” or “crow’s feet”, which can capture the activated degree of AU6. Furthermore, descriptions can also show the inference ability by providing the relationship between current AU and other AUs or emotions, e.g., “indicative of a frown or a scowl”, “associated with strong emotions”. Such descriptions can not only provide more nuances but also improve the interpretability [60, 12] of the model.
@@ 3.2 Instruction-following FABA benchmark
Our REGE score is defined to consider both text generation and image recognition aspects of FABA models. For text generation, we choose to use the ROUGE score, which is a generally used metric in NLG by comparing the overlap of n-grams, word sequences, and word pairs between the generated texts and the reference ones. For recognition, we adopt the recognition accuracy for the multi-class performance of facial emotion recognition, and the F1 score for the multi-label performance of AU recognition. Denote the recognition performance as $S_{re}$ and the generation metric as $S_{ge}$, our REGE metric is computed by taking their sum: $S_{rege}=S_{re}+S_{ge}$.
@@ 4 EmoLA: An Instruction-tuned MLLM for FABA
After obtaining the visual embedding tokens $H_{v}$, facial prior token $H_{p}$ and the language embedding tokens $H_{q}$, we concatenate them together and feed them into the LLM decoder. Here we use Vicuna [15] as the LLM decoder. As shown in Fig. 7, the visual encoder, prior encoder, word embedding and the LLM decoder are frozen. Except for the prior encoder $f_{p}(\cdot)$, the initial weights of the other modules come from a pretrained MLLM. In this paper, we utilize the LLaVA-1.5 [70] as the backbone model. In order to train efficiently without tuning the entire MLLM, we propose to tune an extra LoRA module $h_{\phi}(\cdot)$, a visual projector $h_{\gamma}(\cdot)$, and the prior projector $g_{\theta}(\cdot)$. Compared to finetuning the entire LLM, tuning LoRA reduces both memory and computation costs during training without inducing additional expenses for inference. Our experiments demonstrate the effectiveness of such a design in Tab. 7. As a result, the overall parameters to be optimized are $\boldsymbol{\Theta}=\{\theta,\gamma,\phi\}$. We optimize these parameters following the auto-regressive way, and the likelihood of the generated description $X_{A}$ conditioned on image $X_{V}$, prior facial feature $Z_{P}$ and instructions $X_{Q}$ is given by:
$p\left(X_{A}|X_{V},Z_{P},X_{Q}\right)=\prod_{i=1}^{L}{p_{\boldsymbol{\Theta}}\left(x_{i}|X_{V},Z_{P},X_{Q},X_{A,<i}\right)},$ (2)
@@ 5.2 Comparison on FABA-Bench
We compare with current MLLMs, i.e., LLaVA-1.5 [70], MiniGPT4-V2 [10], Shikra [11], and mPLUG-Owl2 [131], on our FABA-Bench in Tab. 7. These baselines mainly employ Vicuna [15] as LLM decoder and incorporate instruction tuning. We reproduce all the baselines using their open-sourced codes. For fair comparison, all the models are trained for 1 epoch and train on two tasks individually. Except for our EmoLA, all the other baselines are finetuned based on the pretrained MLLMs. More details about baselines can refer to the Appendix.
It is noteworthy that EmoLA achieves the best results in two tasks with fewer tuning parameters compared to other MLLMs. EmoLA finetunes merely 10% of the parameters compared to LLaVA-1.5, yet it achieves better performance on FABA-Bench, which can be attributed to two aspects. For one thing, EmoLA only finetunes LoRA which is much more efficient than tuning the entire LLM decoder. For another, EmoLA incorporates a facial prior expert to extract the facial structure knowledge. This feature compensates for the information overlooked by the visual encoder in FABA tasks. Moreover, we can also see that due to the strong capability of LLM, the language generation ability of Shikra, LLaVA-1.5 and EmoLA is comparable, making it unreasonable to evaluate these models merely based on the NLG metrics. Our metric, REGE, accounts for both the recognition and generation abilities of MLLMs in FABA, offering a more comprehensive evaluation of their performance.
5.3 Ablation study
Table 8: Influence of prior token. $H_{v}$ and $H_{p}$ indicate visual tokens and the facial prior token, respectively.
@@ 5.3 Ablation study
64.5 31.7 56.3 35.2
Compared to merely inputting the prior token $H_{p}$, incorporating visual tokens $H_{v}$ will enhance the recognition performance since these visual tokens carry more detailed face information. Incorporating both visual tokens and prior tokens leads to further improvements with eligible performance dropping in text generation. This suggests that the face prior token supplies complementary information of visual tokens. We suppose this additional information includes facial structure details introduced by landmark facial prior extractor.
Figure 8: An example of EmoLA’s capability on FABA tasks.
Table 9: Influence of the tuning strategy. $h_{\phi}+h_{\gamma}$ and $g_{\theta}$ indicate the LoRA module with visual projector and the prior projector, respectively.
@@
64.5 31.7 56.3 35.2
We also tried exclusively tuning LoRA with visual projector $h_{\phi}+h_{\gamma}$ while keeping other parameters frozen. This is different from merely inputting the visual tokens $H_{v}$ since it also includes a prior token $H_{p}$ generated by the randomly initialized prior projector $g_{\theta}$. The results indicate that tuning $h_{\phi}+h_{\gamma}$ significantly improves the performance by better aligning the output of the LLM. It can also be observed that the performance slightly exceeds that achieved using only the visual tokens, which also emphasizes the value of the prior token. Again, the efficacy is further augmented by simultaneously fine-tuning both $h_{\theta}+h_{\gamma}$ and the prior projector $g_{\theta}$.
@@ 6 Conclusion
Acknowledgement: Yifan Li, Wentao Bao, and Yu Kong are partially supported by NSF Awards 1949694 and 2040209. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NSF. We also would like to express our deepest gratitude to Dr. Junwen Chen for her invaluable support and contribution to this research.
@@ 0.A.1 Limitations
This work also has its limitations. Firstly, we haven’t tried other face-prior feature extractors except for landmark features. We leave this as an exploring research direction for future work. Secondly, our annotations from the training set also include noise induced by GPT-4V’s hallucinations [111, 112] (see Sec. 0.C.3), which may introduce some bias for models. We think our benchmark can be regarded as the learning from noisy descriptions task. Thirdly, the metric we proposed for evaluating the expression recognition ability may not reflect the nuanced expressions, which could be further improved.
@@ 0.C.3 The accuracy estimation of training annotations
Table 11: Estimated accuracy of Emotion annotations in FABA-Instruct.
Total Correct Acc.
200 182 91.0
Table 12: Estimated F1 score of AU annotations in FABA-Instruct.
AU1 AU2 AU4 AU5 AU6 AU7 AU9 AU10 AU12 AU15
92.2 95.6 84.2 90.7 82.6 65.9 61.3 58.1 84.3 70.6
AU17 AU20 AU23 AU24 AU25 AU26 AU27 AU43 Avg.
66.7 55.2 79.1 79.5 84.9 83.9 78.8 57.1 76.1
@@ 0.D.2 Baseline details of FABA-Bench
EmoLA (single task) 64.5 31.7 96.2 56.3 35.2 91.5
@@ 0.D.3 Multi-task performance on FABA-Bench
EmoLA (prior token before visual tokens) 63.5 32.1 95.6 55.4 34.3 89.7
EmoLA (prior token after visual tokens) 64.5 31.7 96.2 56.3 35.2 91.5